# The temperature and metallicity distributions of the ICM: insights with TNG-Cluster for XRISM-like observations

### v1

@@ 3.1 The complexity of the thermal and chemical structure of cluster centres as predicted by TNG-Cluster
Temperature. Starting with the temperature distribution of the gas in the centre of our clusters (Figure 2, top), we find that, according to TNG-Cluster, the majority of the ICM mass is concentrated around a single temperature. The location of the peak exhibits, as expected, a strong correlation with halo mass, with the central ICM of more massive clusters being, on average, hotter. The width of the temperature distributions varies significantly across clusters, even between systems with similar halo masses. A subset of systems, independent of halo mass, appear to have extended tails of gas, extending to sub-keV temperatures. Considering that we are looking at the centrally-located gas, it is very likely that those tails are associated with the cooler inter-stellar (ISM) or circum-galactic (CGM) media of the brightest cluster galaxy (BCG; see also Rohr et al., 2024). A few clusters also exhibit very high, super-virial temperature tails, which, given the SMBH feedback model implemented in TNG-Cluster, may be driven by central energy injections (Truong et al., 2024).
@@
Both the width and shape of the temperature distribution are relevant aspects for our analysis. So, what is the functional form of the temperature of the ICM at the centres of clusters favoured by TNG-Cluster? We fit the temperature distributions with both the Normal and Log-Normal distributions. In Figure 3, we compare the average and best-fit temperature distributions of clusters in halo mass bins, for the intrinsic, mass-weighted gas in the 3D cluster centres (top) and for the emission-weighted gas in a XRISM-like FoV (bottom). The corresponding best-fit values of the Log-Normal distribution for each cluster are shown in Figure 4.
We see that the Log-Normal distributions are better descriptors of the temperature. They give a more accurate representation of the overall temperature distribution, even though the cooler sub-keV gas is systematically underestimated. On the other hand, the Normal distribution naturally accounts for the cooler sub-keV gas, but systematically underestimates the width of the underlying distribution towards higher temperatures, especially in more massive halos, and totally misses the peak of the distributions. A drawback of the Log-Normal distribution is that, when considering the emission-weighted temperature of the gas along the line of sight, the log-Normal model tends to overestimate the hotter tails. Based on Figure 3, the main driver behind this broader width is the enhanced contribution of the sub-keV gas. Because of our weighting scheme, denser gas cells, such as those associated with the ISM or CGM of the BCG, have a stronger contribution on the final distribution and are therefore driving the best-fit width. On the contrary, based on its construction, the Normal best-fit remains fairly unaffected by this enhancement, but it still appears to underestimate the amount of hotter gas present in the emission-weighted temperature distribution.
@@
Overall, our analysis suggests that, while both models are reasonable approximations of the underlying thermal properties of the gas and certainly preferable to single or double temperature models, the Log-Normal functional form is a more accurate representation of the shape of the temperature distribution of the central ICM predicted by TNG-Cluster, consistent with previous works (e.g. Kawahara et al., 2007; Zhuravleva et al., 2013; Khedekar et al., 2013). In Figure 4 we hence provide the best-fit central temperature and standard deviation for the Log-Normal fits, colour coded by the total mass of the cluster. Whereas there is a clear, and known, trend between average central ICM temperature and total cluster mass (spanning between a few keV to more than 10 keV for the studied clusters with log10M at ), the ICM in each cluster exhibit a diverse range of temperatures, irrespective of total halo mass. According to TNG-Cluster, the Log-Normal width of the central ICM temperature distribution can vary between 0.2 and 10 keV. We note that systems with broader distribution widths have extended sub-keV tails, so it is possible that their best-fit widths are biased high due to the ISM or CGM of the BCG.
@@
Upon inspection of all individual systems, the distributions can be broadly described as uni-modal, with the majority of the mass centred around a central peak for the largest majority of the clusters. In fact, we confirm that only a handful of metal-rich clusters exhibit secondary peaks at higher Fe abundances. However, unlike the temperature distributions, the peaks of the Fe abundance distribution show little to no correlation with halo mass: namely, clusters within the same halo mass bins can have significantly different amounts of Fe in their centres. We argue that this is related to their assembly and SMBH feedback histories, and we will expand on this in future works. Indeed, with the TNG model, simulated clusters have been shown to exhibit different central metallicity depending on their cool-core status (Vogelsberger et al., 2018), akin to what suggested by observations. Finally, similar to the tails of cooler gas found in the temperature distributions, a lot of systems tend to have extended tails towards super-solar Fe abundances from gas cells that appear to be associated with the ISM or CGM of the BCG.
@@ 3.2 Inferring the ICM temperature distribution
bvapec | 352 | 100 |
double bvapec | 165 | 47 |
bvgadem | 190 | 54 |
bvlognorm | 213 | 61 |
[…]
Firstly, before quantifying the inference errors, we comment on how frequently it is even possible to successfully fit the X-ray spectra of XRISM/Resolve-like observations of the centres of clusters simulated in TNG-Cluster as a function of underlying spectral-modelling choices. Throughout, we adopt the mocking procedure described in Section 2.2 and the fitting analysis and operational definitions of “accepted” and “rejected” fits of Section 2.3.
Based on the success rates reported in Table 1, the bvlognorm model is the best performer, followed by the bvgadem model and the double bvapec, in that order. The bvapec model always provides a valid best fit: however, by assuming a delta function for the temperature distribution, it fails a priori to provide a realistic depiction of the ICM’s thermal properties. For the other models, the success rate can be taken as a measure for the compatibility between the intrinsic and the assumed temperature distributions. Both the bvlognorm and bvgadem models better describe unimodal albeit broad distributions, which is a reasonable approximation for the majority of TNG-Cluster systems. On the other hand, the double bvapec model is performing less well due to a lack of any visible bimodality, i.e. with two similarly-prominent peaks, in the simulated clusters.
@@
Based on Figure 5, top panels, all models tend to under-predict the mass-weighted mean temperature of the X-ray emitting gas that is actually located in the centre of each cluster. Whereas the bias is systematic, it is in fact relatively small. All models, on average i.e. across the TNG-Cluster sample, underestimate the temperature by 0.8 keV (namely 10 per cent), but with a strong dependence on cluster mass. Namely, the best-fit temperature can underestimate the true one by about 2 keV ( 20 per cent) at the high-mass end, while being a reasonable approximation for lower-mass clusters. This result underscores the difficulty of assessing the true physical state of central ICM gas. In turn, assuming that TNG-Cluster returns reasonably realistic systems, the top panels of Figure 5 can provide guidelines to correct for this observational-analysis bias. This is the case on average, whereas for individual clusters the inference error can be much larger – we expand on outlier cases later.
In terms of emission-weighted quantities and comparing to the true properties of all the gas along the line of sight of a XRISM-like central pointing (Figure 5, bottom panels), the situation is better, as expected. All models, but to a lesser degree the double bvapec one, produce a best-fit temperature that is indicative of the emission-weighted mean temperature of the gas for a large majority of the studied clusters. The good accuracy of this inference appears independent of halo mass, but the scatter seems to increase somewhat towards the low-mass end. For the double bvapec model, in particular, and for both the top and bottom panels, the inference errors span the largest ranges, with a sizeable fraction of clusters whose observational-based (primary) temperature is substantially higher or lower than the true one. Moreover, the errors on the best-fit values for the bvapec model are on average much larger than in any other case (red error-bars).
@@
In Figure 6, we quantify the residuals for both the mass-weighted temperature distribution of the central gas (top) and the emission-weighted gas in a XRISM-like FoV (bottom). For both cases, the bvgadem model tends to overestimate the width of the distribution, with an average residual of 1.5 (-1.9,+1.7) and 1.5 (-1.6,+1.6), for top and bottom cases, respectively. On the other hand, the bvlognorm model has smaller on average residuals for the central mass-weighted gas (0.6 (-1.2,+2.3)) but tends to underestimate the width of the emission-weighted gas in the XRISM FoV by an average of -2.1 (-0.01,+3.6). We note that the average value of the bvlognorm residuals in the latter case is biased low by a handful of systems for which the model severely underestimates the best-fit width of the emission-weighted temperature distribution.
However, the strong outliers may not arise from the spectral fitting itself, but rather from the inflated value of the best-fit Log-Normal width to the intrinsic gas distribution. As we have shown in Figure 3, the sub-keV gas drives the width of the best-fit Log-Normal distribution towards systematically higher values, due to our weighting scheme, biasing the bvlognorm inferences low.
@@ 3.3 Inferring the ICM Fe abundance
In Figure 7 we show the comparison between the inferred Fe abundance from all spectral-emission models compared to the mass-weighted mean Fe abundance of the ICM in the actual 3D cluster centre (top) and the emission-weighted mean Fe abundance of gas from a central pointing with a XRISM-like FoV (bottom). As the top panel of Figure 7 shows, the inference on the Fe abundance is systematically different of the actual amount of Fe in cluster centres. Namely, all models underestimate the central Fe abundance by 0.11 Solar (21 per cent). The offset varies with halo mass, being larger towards the low- and high-mass end of our mass range. However, the inference error is at least partially driven by the subset of clusters that, despite being actually metal rich, are found to be significantly metal poorer when their X-ray spectra are analysed, irrespective of the adopted spectral-emission model. This discrepancy disappears if we account for emission (bottom panels), suggesting it is possibly driven by less X-ray bright gas. Interestingly, even though the rejected fits do not occupy a special region in the parameter space, as in the temperature case, a disproportional amount of clusters appear to belong to this subset.
Similar to the case of the emission-weighted mean temperature, all models consistently trace the emission-weighted Fe abundance, or at least they trace more accurately the emission-weighted than the mass-weighted Fe abundance, with no significant halo-mass dependence. Only exception to this is the bvlognorm: the latter systematically underestimates the emission-weighted Fe abundance along the LoS, and it is in fact more sensitive to the mass-weighted value.
@@ 4 Discussion
While the inferences from the bvapec model are comparable to (if not more accurate on average than) the rest, the model, by design, ignores the presence of multi-phase gas. Attempting to rectify this with the inclusion of another single-temperature component, in the form of the double bvapec model, does not improve the accuracy of the inferences on average (i.e. across the studied cluster sample) and actually returns a much larger fraction of clusters with very large inference errors on the ICM effective temperature, with biases in both directions.
@@
As we have previously shown, all the spectral-emission models used in this work are capable of inferring with good accuracy the emission-weighted summary statistics of the cluster gas from a central pointing with a XRISM-like observation. Yet, for a large number of clusters, they fail to return an accepted best fit of the underlying X-ray spectrum. Moreover, as mentioned above and by comparing the filled and empty markers of Figure 5, Figure 6 and Figure 7, it would appear that the rejection of the X-ray spectra fits (at least based on our criteria, see subsubsection 2.3.1) does not translate in much or overall worse inference errors. This can be appreciated via the quantitative summaries provided in Table 2 and Table 3. In the following, we hence provide further considerations as to why the quality of the fit may still be relevant when choosing which spectral-emission model to adopt to analyse the data.
@@
In Figure 8 we showcase the emission-weighted distributions of the ICM properties of a few example individual systems in comparison to their best-fit description, for all the spectral-emission models studied in this work. Each cluster is represented by two panels, one for temperature and one for Fe abundance, respectively, on top of each other. We show a total of 12 example clusters, selected based on the performance of the bvlognorm model to infer, respectively, effective temperature (top two rows), Fe abundance (bottom two rows) and temperature width (middle two rows). We focus on the bvlognorm because it is the best descriptor of the TNG-Cluster temperature distributions (Figure 3 and Figure 4).
@@
From left to right, we visualize the intrinsic and inferred gas property distributions for different example cases. In the leftmost column, we show systems with the minimum absolute difference between the true and inferred value of the relevant property. We also showcase instances of an overestimation and underestimation, located at the 95 and 5 percentile of the sorted difference distribution, respectively (middle columns). Finally, we also show examples of rejected fits (rightmost column), so that we can better understand the intrinsic properties of the gas that, despite an accurate inference, led to model rejection.
@@
Starting with the accepted spectral fits, we can understand how each model attempts to infer the input temperature distribution. The double bvapec model (red lines) systematically infers temperatures that are not present or not associated with any minor peaks in the temperature distribution. Similarly, in many cases it fails to identify the most prominent peak, by inferring two temperatures around it. This suggests that the model, in an attempt to approximate a distribution of temperatures, does not provide an accurate representation of the actual distribution but an attempt of residual minimisation.
@@
The bvgadem model (blue lines) can provide a good approximation of distributions extended towards higher temperatures, but systematically assumes a much larger amount of cooler gas compared to what is present. Finally, many cases show how the bvlognorm model provides a fairly accurate representation of the width of the temperature distributions. However, it tends to fail when the intrinsic thermal structure deviates strongly from a Log-Normal distribution in the form of extended tails towards higher- or lower-temperature gas.
@@
Regarding the Fe abundance distributions, all spectral-emission models make overly simplistic approximations. We see a range in the capability of all models to identify the Fe abundance of the majority of the gas, with the bvlognorm model inferring lower values compared to the other models. This behaviour is consistent with our results above, where the models appear more sensitive to the mean values.
@@
Another important aspect regarding the inference of each model are the conditions that led to its rejection. As we can see in Figure 8 and as previously indicated in Figure 5 and Figure 7, the rejected fits of all models do infer a fairly accurate value of the mean temperature and Fe abundance emission-weighted distribution. For bvgadem and bvlognorm, the main and most frequent factor that leads to the rejection of the model is the width of the distribution: in some rejected cases, the inferred width is simply too large, infinity; in other rejected cases, the model infers a very narrow temperature distribution, i.e. with a poor constraint on the width itself.

### v2

@@ 3.1 The complexity of the thermal and chemical structure of cluster centres as predicted by TNG-Cluster
Temperature. Starting with the temperature distribution of the gas in the centre of our clusters (Figure 2, top), we find that, according to TNG-Cluster, the majority of the ICM mass is concentrated around a single temperature, the virial temperature of the cluster. The location of the peak exhibits, as expected, a strong correlation with halo mass, with the central ICM of more massive clusters being, on average, hotter. The width of the temperature distributions varies significantly across clusters, even between systems with similar halo masses. A subset of systems, independent of halo mass, appear to have extended tails of gas, extending to sub-keV temperatures. Considering that we are looking at the centrally-located gas, it is very likely that those tails are associated with the cooler inter-stellar (ISM) or circum-galactic (CGM) media of the brightest cluster galaxy (BCG; see also Rohr et al., 2024). A few clusters also exhibit very high, super-virial temperature tails, which, given the SMBH feedback model implemented in TNG-Cluster, may be driven by central energy injections (Truong et al., 2024). Overall, by visual inspection of the temperature distribution of each individual simulated cluster, we identify that at least 20 per cent of them (i.e. about 70 clusters of the 352) exhibit ICM in the central regions extending across more than one order of magnitude in temperature (in units of keV), i.e. with extended tails either towards cooler or super-virial temperatures.
@@
Both the width and shape of the temperature distribution are relevant aspects for our analysis. So, what is the functional form of the temperature of the ICM at the centres of clusters favoured by TNG-Cluster?
We fit the temperature distributions in the 0.1-50 keV range with both the Normal and Log-Normal distributions. In Figure 3, we compare the geometric mean and best-fit temperature distributions of clusters in halo mass bins, for the intrinsic, mass-weighted gas in the 3D cluster centres (top) and for the emission-weighted gas in a XRISM-like FoV (bottom). We note that the average curves are not representative of each individual cluster’s best fit, but rather are meant to highlight possible systematic trends with cluster mass. The corresponding best-fit values of the Log-Normal distribution for each cluster are shown in Figure 4. The quality of each fit on a cluster-by-cluster basis is provided by the residuals in Figure 16, and it is summarized by the median across the entire cluster sample annotated as text in Figure 3.
As Figure 3 suggests, the temperature distribution cannot be perfectly described by either a Normal or a Log-Normal distribution in all systems. Both functional forms appear to be a reasonably good descriptor of the bulk of the temperature distribution around the virial temperature peaks. However, both models introduce biases primarily due to the presence of extended tails of gas in either direction, i.e. in those clusters where such tails are present. Both for the mass-weighted, centrally located gas as well as for the emission-weighted gas along the line of sight, the best-fit Normal distribution may underestimate to some small degree the amount of virial and hotter gas while typically systematically overestimating the amount of cooler gas, compared to the mode, present in the ICM.444Whereas the biases in the fits appear more pronounced in the most massive halos, this is an artifact of the averaging across multiple single-cluster distributions.
On the other hand, the Log-Normal model appears to be totally insensitive to the extended tails of sub-keV gas, which it fails to capture. This effect is also reflected in the , where, for both gas selections, the model tends to under-fit the true distribution of temperatures. Amid this overall picture, the cluster-to-cluster variation remains significant in all cases, as can be appreciated in Figure 16. From the visual inspection of the fits to each cluster temperature distribution, the Log-Normal model also tends to grossly underestimate the amount of gas at the hottest-end of the temperature distribution (in about 100 clusters of 352).
Overall, our analysis suggests that, a priori, neither functional form provides a perfectly accurate representation of the underlying thermal properties of the gas for all clusters. While both the Normal and Log-normal distributions are certainly more informative than a single or double delta function, they can not fully capture the tails towards cooler and hotter gas present in the intrinsic temperature distributions of a non-negligible fraction of clusters. It is evident that more sophisticated models are needed to fully capture the thermal structure of the ICM, but a more in depth analysis and the development of such models is outside the scope of this paper.
@@
In Figure 4 we provide the best-fit central temperature and standard deviation for the Log-Normal fits, colour coded by the total mass of the cluster. Whereas there is a clear, and known, trend between average central ICM temperature and total cluster mass (spanning from a few keV to more than 10 keV for the studied clusters with log at ), the ICM in each cluster exhibits a diverse range of temperatures, irrespective of total mass. According to TNG-Cluster, the Log-Normal width of the central ICM temperature distribution varies between 0.4 and 10 keV.
@@
Upon inspection of all individual systems, the distributions can be broadly described as uni-modal, with the majority of the mass centred around a central peak for the largest majority of the clusters. In fact, we confirm that only a handful of metal-rich clusters exhibit secondary peaks at higher Fe abundances. However, unlike the temperature distributions, the peaks of the Fe abundance distribution show little to no correlation with halo mass: namely, clusters within the same halo mass bins can have significantly different amounts of Fe in their centres. We argue that this is related to their assembly and SMBH feedback histories, and we will expand on this in future works. Indeed, with the TNG model, simulated clusters have been shown to exhibit different central metallicity depending on their cool-core status (Vogelsberger et al., 2018), akin to what is suggested by observations. Finally, similar to the tails of cooler gas found in the temperature distributions, a lot of systems tend to have extended tails towards super-solar Fe abundances from gas cells that appear to be associated with the ISM or CGM of the BCG.
@@ 3.2 Inferring the ICM temperature distribution
| bvapec | 322 | 91 |
| double bvapec | 154 | 44 |
| bvgadem | 160 | 45 |
| bvlognorm | 199 | 57 |
As shown and discussed above, as well as in Rohr et al. (2024) and Staffehl et al. (2025), the intrinsic temperature distribution of the ICM, as predicted by TNG-Cluster, is multiphase in nature. Therefore, an important aspect for the accuracy of spectral-emission models is their ability to reflect such a multiphaseness in their best fits – we tackle this next.
First, before quantifying the inference errors, we comment on how frequently it is even possible to successfully fit the X-ray spectra of XRISM/Resolve-like observations of the centres of clusters simulated in TNG-Cluster as a function of underlying spectral-modelling choices. Throughout, we adopt the mocking procedure described in Section 2.2 and the fitting analysis and operational definitions of “accepted” and “rejected” fits of Section 2.3.
Based on the success rates of the various models reported in Table 1, the bvlognorm model is the best performer among the multi-temperature ones, followed by the bvgadem model and the double bvapec, in that order. The bvapec model has, by far, the largest number of accepted fits: however, by assuming a delta function for the temperature distribution, it fails a priori to provide a realistic depiction of the ICM’s thermal properties. For the other models, the success rate can be taken as a measure for the compatibility between the intrinsic and the assumed temperature distributions. For example, the bvlognorm model assumes a more accurate depiction of the ICM temperature distribution as a unimodal distribution and remains insensitive to the sub-keV gas, which should have a negligible contribution in a XRISM/Resolve like spectrum. On the other hand, the bvgadem model, which tends to overestimate the contribution of the sub-keV gas, underperforms due to inflated distribution widths. With a similar acceptance rate, the double bvapec model also underperforms since the simulated clusters lack any visible bimodality, i.e. two similarly-prominent peaks.
@@
Based on Figure 5, top panels, all models tend to under-predict the mass-weighted mean temperature of the X-ray emitting gas that is actually located in the centre of each cluster. Whereas the bias is systematic, it is in fact relatively small. All models, on average i.e. across the TNG-Cluster sample, underestimate the temperature by 1.19 keV (an error of 17 per cent), but with a strong dependence on cluster mass. Namely, the best-fit temperature can underestimate the true one by about 1.5 keV (20 per cent) at the high-mass end, while being a reasonable approximation for lower-mass clusters. This result underscores the difficulty of assessing the true physical state of central ICM gas. In turn, assuming that TNG-Cluster returns reasonably realistic systems, the top panels of Figure 5 can provide guidelines to correct for this observational-analysis bias. This is the case on average, whereas for individual clusters the inference error can be much larger – we expand on outlier cases later.
In terms of emission-weighted quantities and comparing to the true properties of all the gas along the line of sight of a XRISM-like central pointing (Figure 5, bottom panels), the situation improves. All models produce a best-fit temperature that is indicative of the emission-weighted mean temperature of the gas for a large majority of the studied clusters, independent of halo mass. For the double bvapec model, in particular, and for both the top and bottom panels, the inference errors span the largest ranges, with a sizeable fraction of clusters whose observational-based (primary) temperature is substantially higher or lower than the true one. Moreover, the errors on the best-fit values for the double bvapec model are on average much larger than in any other case.
@@
In Figure 6, we quantify the residuals for both the mass-weighted temperature distribution of the central gas (top) and the emission-weighted gas in a XRISM-like FoV (bottom). For both cases, the bvgadem model tends to overestimate the width of the distribution, with an average of 1.56 (-1.77,+0.55) and 1.27 (-1.67,+0.56), for top and bottom cases, respectively. On the other hand, the bvlognorm model tends to overestimate the average for the central mass-weighted gas (1.05 (-1.28,+0.70)), but is in a good agreement with the width of the emission-weighted gas in the XRISM FoV by an average of 0.41 (-1.05,+0.85). We note that, for both models, we find inferences that deviate by more than 5 from the true width of the distribution, even for accepted fits. These cases appear to be systems with temperature distributions that have pronounced tails of cooler or super-virial gas.
@@ 3.3 Inferring the ICM Fe abundance
In Figure 7 we show the comparison between the inferred Fe abundance from all spectral-emission models compared to the mass-weighted mean Fe abundance of the ICM in the actual 3D cluster centre (top) and the emission-weighted mean Fe abundance of gas from a central pointing with a XRISM-like FoV (bottom). As the top panel of Figure 7 shows, the inference on the Fe abundance is systematically different of the actual amount of Fe in cluster centres. Namely, all models underestimate the central Fe abundance by about 0.12 Solar (22 per cent). The offset varies with halo mass, being larger towards the low-mass end of our mass range. It is important to point out that such a systematic offset is significantly larger than the statistical errors on the Fe abundance typically quoted in the literature, of the order of 0.01, e.g. for the inferences from XRISM/Resolve observations available in the literature.
However, the inference error is at least partially driven by the subset of clusters that, despite being actually metal rich, are found to be significantly metal poorer when their X-ray spectra are analysed, irrespective of the adopted spectral-emission model. This discrepancy disappears if we account for emission (bottom panels), suggesting it is possibly driven by less X-ray bright gas. Interestingly, even though the rejected fits do not occupy a special region in the parameter space, as in the temperature case, a disproportional amount of clusters appear to belong to this subset.
Similar to the case of the emission-weighted mean temperature, all models consistently trace the emission-weighted Fe abundance, or at least they trace more accurately the emission-weighted than the mass-weighted Fe abundance, with no significant halo-mass dependence.
@@ 4 Discussion
As shown in Figure 5 and Figure 7, the inferences of the bvapec model on the emission-weighted temperature and Fe abundance of the gas along the LoS are comparable to (if not more accurate on average than) the other spectral-emission models. We note that for the temperature, this is achieved despite us not using the so-called spectroscopic temperature weighting from Mazzotta et al. (2004), which was introduced to account for biases on the inference from CCD-resolution mock observations. Similarly, we find a good agreement between the true and inferred Fe abundance, even for haloes with temperatures between 2 and 3 keV, suggesting that our estimates are not subjected to the "inverse Fe-bias" (e.g. Rasia et al., 2008; Simionescu et al., 2009; Gastaldello et al., 2010). We believe that, in both cases, the bvapec model is able to provide a more accurate inference on the true temperature and Fe abundance of the ICM due to the higher spectral resolution, which reveals more lines that can place strong constraints on the model. However, our analysis remains limited to the cluster centres and the quality of the inference varies from target to target, suggesting that a broader examination of this topic, for a large sample of galaxy clusters over larger scales is necessary, but outside the scope of this project.
However, the bvapec model, by design, ignores the presence of multi-phase gas and cannot provide us any insights on the distribution of temperatures present in the ICM. Attempting to rectify this with the inclusion of another single-temperature component, in the form of the double bvapec model, does not improve the accuracy of the inferences on average (i.e. across the studied cluster sample) and actually returns a much larger fraction of clusters with very large inference errors on the ICM effective temperature, with biases in both directions.
@@
As we have previously shown, all the spectral-emission models used in this work are capable of inferring with good accuracy the emission-weighted summary statistics of the cluster gas from a central pointing with a XRISM-like observation. Yet, for a large number of clusters, they fail to return an accepted best fit of the underlying X-ray spectrum. Moreover, as mentioned above and by comparing the filled and empty markers of Figure 5, Figure 6 and Figure 7, it would appear that the rejection of the X-ray spectra fits (at least based on our criteria, see Section 2.3.1) does not translate in much or overall worse inference errors. This can be appreciated via the quantitative summaries provided in Table 2 and Table 3. In the following, we hence provide further considerations as to why the quality of the fit may still be relevant when choosing which spectral-emission model to adopt to analyse the data.
@@
In Figure 8 we showcase the emission-weighted distributions of the ICM properties of a few example individual systems in comparison to their best-fit description, for all the spectral-emission models studied in this work. Each cluster is represented by two panels, one for temperature and one for Fe abundance, respectively, on top of each other. We show a total of 12 example clusters, selected based on the performance of the bvlognorm model to infer, respectively, effective temperature (top two rows), Fe abundance (bottom two rows; continued Figure) and temperature width (middle two rows).
We focus on the bvlognorm because it provides a good descriptor of the TNG-Cluster temperature distributions (subsection 3.1) and because it provides more frequently more accurate best-fit results for the temperature distribution widths (subsection 3.2). However, this discussion can be generalized for all spectral-emission models.
@@
From left to right, we visualize the intrinsic and inferred gas property distributions for different example cases. Similar to Figure 3, we present the probability mass function of the emission-weighted gas distribution along the LoS as a fraction of the EM per bin. In the leftmost column, we show systems with the minimum absolute difference between the true and inferred value of the relevant property. We also showcase instances of an overestimation and underestimation, located at the 95 and 5 percentile of the sorted difference distribution, respectively (middle columns). Finally, we show examples of rejected fits (rightmost column), so that we can better see the intrinsic properties of the gas that, despite an accurate inference, led to model rejection.
@@
Starting with the accepted spectral fits, we can understand how each model attempts to infer the input temperature distribution. The double bvapec model systematically infers temperatures that are not present or not associated with any minor peaks in the temperature distribution. Similarly, in many cases it fails to identify the most prominent peak, by inferring two temperatures around it. This suggests that the model, in an attempt to approximate a bimodal distribution of temperatures, does not provide an accurate representation of the actual distribution, but an attempt of residual minimisation.
@@
The bvgadem model is capable of determining the peak of the temperature distribution, but systematically assumes a much larger amount of cooler gas compared to what is present. Finally, many cases show how the bvlognorm model provides a fairly accurate representation of the width of the temperature distributions. However, it tends to fail when the intrinsic temperature distribution deviates strongly from a Log-Normal distribution in the form of extended tails towards higher- or lower-temperature gas.
@@
Regarding the Fe abundance distributions, all spectral-emission models make overly simplistic approximations. We see a range in the capability of all models to identify the mean Fe abundance, consistent with the trends found above.
@@
Another important aspect regarding the inference of each model are the conditions that led to its rejection. As we can see in Figure 8 and as previously indicated in Figure 5 and Figure 7, even rejected fits of all models are capable of inferring a fairly accurate value of the mean temperature and Fe abundance emission-weighted distribution. For bvgadem and bvlognorm, the main and most frequent factor that leads to the rejection of the model is the width of the distribution: in some rejected cases, the inferred width is simply too large, infinity; in other rejected cases, the model infers a very narrow temperature distribution, i.e. with a poor constraint on the width itself.